The discriminant of a general quadratic equation ax2+bx+c = 0, where a≠0, is given by D = b2−4ac. 

i.   State three possibilities of the discriminant.

ii.   Use the discriminant to determine how many real solutions the equation 2x2 −10x + 20 = 0 have?   


1
Expert's answer
2021-03-07T14:24:21-0500

1)If D>0, we will have 2 different real numbers.

If D=0, we will have 2 equal numbers.

If D<0, we will have complex conjugate numbers

2)

"2x^2-10x+20=0\\
x^2-5x+10=0\\
D=25-40<0"

It means that equation won't have a solution in real numbers